You have taken a long position in a call option on IBM common stock. The option has an exercise price of $144 and IBM's stock currently trades at $148. The option premium is $6 per contract. a. How much of the option premium is due to intrinsic value versus time value? b. What is your net profit on the option if IBM’s stock price increases to $158 at expiration of the option and you exercise the option? c. What is your net profit if IBM’s stock price decreases to $138?
Answer #1

intrinsic value = market price - strike = 148 -144 = 4

profit @ 158 = price-strike-option premium = 158-144-6 = 8

profit @ 138 = min(price-strike,0)-option premium = min(138-144,0)-6 = 0-6 = -6